40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3979     Entry time: Tue Nov 23 18:08:28 2010
Author: Jenne 
Type: Update 
Category: SUS 
Subject: ETMU07: Balanced, standoff glued. ETMU05: Magnets glued to optic 

[Koji, Jenne]

ETMU07 had its wire winched to the correct height, was balanced, standoff glued.  Can be ready for going into the oven tomorrow, if an oven is available.  (One of Bob's ovens has a leak, so he's down an oven, which puts everything behind schedule.  We may not be able to get anything into the oven until Monday).

ETMU05 had magnets glued to the optic.  Hopefully tomorrow we will winch the wire and balance the optic, and glue the standoff, and be ready to go into the oven on Monday.

The spring plungers were sonicated, but have not yet been baked.  I told Daphen that we'd like the optics baked first, so that we can get ETMX in the chamber ASAP, and then the spring plungers as soon as possible so that we can install ETMY and put the OSEMs in.

The updated status table:

StatusTable.png

ELOG V3.1.3-